OpenStudy (anonymous):

what is the square root of 52?? Ill give first one to answera medal!!

OpenStudy (btaylor):

To find it, write out the factors: \(\sqrt{52}=\sqrt{2 \times 2 \times 13} \). Since there are two 2s inside the square root, you can move them to the outside to get \(2 \sqrt{13}\).

OpenStudy (anonymous):

so thats the answer??

OpenStudy (btaylor):

yep. If you need the approximate decimal answer you can punch square root of 52 into your calculator. It comes out to 7.21110255...

OpenStudy (anonymous):

ooooohhhhh... thanks!!!!!

OpenStudy (btaylor):

you're welcome.